### v1

@@ 6 Concluding remarks
We have argued that blockchain-based smart contracts enable credible formulation of reaction functions. Such an approach is particularly relevant in case of investment games in which players can deposit funds and condition their investment on the investment levels of the others. The outcome that is implemented is then governed by a self-executable coordinating contract. As such, reaction functions can help non-cooperative players act out centralized coordination in order to resolve issues of trust (as in weakest-link games) and free-riding (as in public-good games).
Compared to the approaches in Tennenholtz (2004) and Kalai et al. (2010), we submit that the reaction functions bring automated play a step closer to practical application. For example, to obtain collaboration as a program-equilibrium outcome in the Prisoner’s dilemma, Tennenholtz (2004) suggests a program of the form “if my program is identical to my opponent’s program, then the program collaborates; otherwise, it defects”. Conditioning your action on the program of the opponent is fragile, as the smallest difference in programs yields defection. Even though players may be willing to collaborate, programs can be formulated in countless ways and there is a risk that a collaborative outcome would never prevail in practice. For instance, imagine that the other player is naïve and chooses a program that always collaborates—as the programs differ, this will not lead to collaboration. In contrast, our reaction functions specify an action for any finite action profile of the others. This may be more robust as illustrated in the cases of weakest-link and public-good games. For example, if players exhibit conditional collaboration in a weakest-link game, we will not automatically resort to the worst-case outcome if some players deviate from the best-reply reaction function.
We end on some points for future research. One is to explore reaction functions in games with incomplete information. For instance, the idea of “playing safe” seems appealing even if, say, you do not know the preferences of the other players. A first step in that direction is Theorem 5.4, which shows that the best-reply reaction is weakly dominant against monotone play in high-risk weakest-link games. As weak dominance is a statement only on the player’s own payoffs, knowledge of the others’ payoffs is not important. One can ask more generally, what is required for $\textit{BR}_{i}$ to be weakly dominant against monotone play? Theorem 5.4 then points to some conditions: (a) payoff $u_{i}(a)$ is non-decreasing in $a_{-i}\in A_{-i}$, (b) $\textit{BR}_{i}$ is monotone (to ensure fixed points against monotone play), and (c), for each outcome $a\in A$,
@@
There are many games, different from high-risk weakest-link games, for which this applies. The common theme is that (a) $i$ is not harmed by $j$’s actions, that (b) there are some complementarities between player contributions, and (c) that there is a “target” that the player wants to match but not exceed. This target could for instance be derived from Kantian fairness reasoning (e.g. Roemer, 2010; Alger and Weibull, 2013) combined with a strong disutility from being “exploited” (investing more than what is “right” given the investments of the others to the others’ benefit). Given the prominent position of the best-reply reaction function, which is simple and intuitive, it would be interesting to map out exactly when it is also optimal.
@@
Consider a potential deviation $R^{\prime}_{i}\in\mathcal{R}_{i}$. For each $a^{\prime}\in\mathcal{E}(R^{\prime}_{i},R_{-i})$ that differs from $a^{*}$, $a^{\prime}_{-i}$ minimizes $u_{i}$ given $a^{\prime}_{i}$. That is,
@@
Part II: For contradiction, say $\tilde{R}\in\mathcal{R}$ supports $\tilde{a}\in A$ in reaction-function equilibrium yet $\underaccent{\bar}{v}_{i}>u_{i}(\tilde{a})$ for some player $i$. Let $a^{\prime}_{i}\in A_{i}$ be $i$’s maxmin strategy,
@@
That is, $i$ is influenced only by their “neighbors” and takes action $1$ if the player before them chooses 00 whereas the one after them doesn’t. Figure 1 illustrates this case.
[…]
@@
As $R_{i}(0,\dots,0)=0$, the outcome $(0,\dots,0)$ is a fixed point. Say player $i$ is unhappy at outcome $a$ if $R_{i}(a_{-i})\neq a_{i}$. We will argue that there are at least two unhappy players at each $a\neq(0,\dots,0)$. For instance, $R_{i}(1,\dots,1)=0$, so all players are unhappy at $(1,\dots,1)$. We restrict to the most challenging cases, namely $a\in\{0,1\}^{n}$. It remains therefore to check outcomes $a$ that include at least one 00 and at least one $1$.
Suppose for contradiction that there is at most one unhappy player at $a$. By construction, $a$ contains at least one sequence of consecutive $1$’s (possibly a single $1$) followed by a 00. The last player $m$ of the sequence is unhappy as $a_{m}=1$ and $a_{m+1}=0$. Therefore, $a$ can contain at most one such sequence (else there would be at least two unhappy players). Moreover, each player $i$ not the first of the sequence is unhappy as $a_{i-1}=a_{i}=1$. Hence, the sequence contains at most two players. There are two options: if $a_{i}=1$ and $a_{j}=0$ otherwise, then $i-1$ and $i$ are unhappy; if $a_{i}=a_{i+1}=1$ and $s_{j}=0$ otherwise, then $i-1$ and $i+1$ are unhappy. (In the second case, that $i-1$ is unhappy hinges on $a_{i-2}=0$. As then $a_{i-2}\neq a_{i+1}$, this uses that $n\geq 4$.) Hence, at all outcomes $a\neq(0,\dots,0)$ there are at least two unhappy players.
@@
If a safe outcome $\underaccent{\bar}{a}$ of the strategic game is a Nash equilibrium, then there is a safe reaction-function equilibrium. In general, if a safe reaction-function equilibrium exists, then there exists a Pareto-efficient safe reaction-function equilibrium.
@@
Part I: Let $R^{*}_{i}\in\mathcal{M}_{i}$ be norm-proof. Fix an amount $\alpha\in\{1,\dots,H\}$ and let $R\in\mathcal{N}$ be such that, for a generic player $i$, $R_{i}(a_{-i})=\alpha$ if $\min_{j\neq i}a_{j}\geq\alpha$ and otherwise $R_{i}(a_{-i})=0$. Then $(\alpha,\dots,\alpha)\in\mathcal{E}(R)$ and $U_{i}(R^{*}_{i},R_{-i})\geq U_{i}(R)$ requires $R^{*}_{i}(\alpha,\dots,\alpha)=\alpha$. As this applies for each $\alpha\in\{1,\dots,H\}$, we conclude that $R^{*}_{i}$ must coordinate if all others do.
Let now $a_{-i}\in A_{-i}$ be such that $\min_{j\neq i}a_{j}<\max_{j\neq i}a_{j}\equiv\alpha$. For contradiction, suppose that $R^{*}_{i}(a_{-i})\geq\alpha$. Then $(\alpha,\dots,\alpha)\geqq a_{-i}$. As $R^{*}_{i}$ is monotone, $R^{*}_{i}(\alpha,\dots,\alpha)\geq R_{i}^{*}(a_{-i})>\alpha$. But $R^{*}_{i}(\alpha,\dots,\alpha)=\alpha$, which is a contradiction. Furthermore, as $R^{*}_{i}$ is monotone, $R^{*}_{i}(a_{-i})\geq R^{*}_{i}(\min_{j\neq i}a_{j},\dots,\min_{j\neq i}a_{j})=\min_{j\neq i}a_{j}$. Hence, $R^{*}_{i}$ exhibits conditional collaboration.
Part II: Let $R^{*}_{i}$ exhibit conditional collaboration. For contradiction, suppose $R^{*}_{i}$ is not norm-proof, so there is $R\in\mathcal{N}$ for which $U_{i}(R^{*}_{i},R_{-i})<U_{i}(R)$. We claim that each $a\in\mathcal{E}(R)$ is coordinated: Without loss, say $a\in\mathcal{E}(R)$ is ordered such that $a_{1}\leq\dots\leq a_{n}$. For contradiction, suppose $a_{1}<a_{n}$. Then $a_{-1}\geqq a_{-n}$. As $R_{i}$ is monotone and $a\in\mathcal{E}(R)$, $a_{1}=R_{1}(a_{-1})\geq R_{n}(a_{-n})=a_{n}$, which is a contradiction. Hence, $U_{i}(R)=u_{i}(\alpha,\dots,\alpha)$ for some $\alpha\in\{1,\dots,H\}$. But $R^{*}_{i}(\alpha,\dots,\alpha)=\alpha$, so $U_{i}(R^{*}_{i},R_{-i})\geq u_{i}(\alpha,\dots,\alpha)=U_{i}(R)$. This is a contradiction. ∎
@@
For symmetric investment games, all profiles consisting only of reaction functions that exhibit conditional collaboration support the high-investment outcome $(H,\dots,H)$ in reaction-function equilibrium.
@@
Let $R\in\mathcal{R}$ be such a profile. Then $\mathcal{E}(R)=\{(a_{i},\dots,a_{i})\in A\mid a_{i}\in A_{i}\}$ includes, in particular, the desired outcome $(H,\dots,H)$. Consider an arbitrary deviation $R^{\prime}_{i}\in\mathcal{R}_{i}$ and associated fixed point $a\in\mathcal{E}(R^{\prime}_{i},R_{-i})$. There are two cases. Either $a$ is coordinated, $a=(a_{i},\dots,a_{i})$, and we have $u_{i}(a)\leq u_{i}(H,\dots,H)$ as coordinating on a higher action is better. Otherwise, $a$ is not coordinated. By conditional collaboration, for each player $j\neq i$, $a_{j}=R_{j}(a_{-j})<\max_{k\neq j}a_{k}$. But then we must have $a_{i}>a_{j}$ for each $j\neq i$. As $u_{i}$ is non-decreasing in $a_{j}$ for $j\neq i$, $u_{i}(a)\leq u_{i}(a_{i},\dots,a_{i})$. As coordinating on higher action is better, $u_{i}(a_{i},\dots,a_{i})\leq u_{i}(H,\dots,H)$. In both cases, we conclude that $U_{i}(R^{\prime}_{i},R_{-i})\leq U_{i}(R)$, so there is no beneficial deviation and $R$ is a reaction-function equilibrium. ∎
Theorem 5.3.
For weakest-link games, reaction function $R_{i}\in\mathcal{R}_{i}$ is payoff consistent and exhibits conditional collaboration if and only if $R_{i}=\textit{BR}_{i}$.
Proof.
Let $R_{i}\in\mathcal{R}_{i}$ be payoff consistent and exhibit conditional collaboration. Consider an arbitrary $a\in A$ and let $m=\min_{j\neq i}a_{j}$. By payoff consistency, $R_{i}(a_{-i})=R_{i}(m,\dots,m)$. By conditional collaboration, $R_{i}(m,\dots,m)=m$. Hence, $R_{i}(a_{-i})=\min_{j\neq i}a_{j}=\textit{BR}_{i}(a_{-i})$. ∎
@@
For public-good games, reaction function $R_{i}\in\mathcal{R}_{i}$ is monotone, payoff consistent, and exhibits conditional collaboration if and only if, for each $a_{-i}\in A_{-i}$,
$R_{i}(a_{-i})=R^{*}_{i}(a_{-i})=\left\lfloor\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rfloor.$
@@
Let $R_{i}\in\mathcal{R}_{i}$ be monotone, payoff consistent, and exhibit conditional collaboration. Consider an arbitrary $a\in A$ and let
@@
We can “flatten” $a$ to $a^{\prime}\in A$ with $a^{\prime}_{-i}=(m,\dots,m,m+1,\dots,m+1)$ and $\sum_{j\neq i}a^{\prime}_{j}=\sum_{j\neq i}a_{j}$. By payoff consistency, $R_{i}(a_{-i})=R_{i}(a^{\prime}_{-i})$. If $a^{\prime}_{-i}=(m,\dots,m)$, then conditional collaboration gives $R_{i}(a^{\prime}_{-i})=m$ and thus $R_{i}(a_{-i})=m$, as desired. If not, so $a^{\prime}_{-i}\neq(m,\dots,m)$, then $a^{\prime}_{-i}$ contains at least one $m$ and at least one $m+1$. By monotonicity, $R_{i}(m,\dots,m)\leq R_{i}(a^{\prime}_{-i})\leq R_{i}(m+1,\dots,m+1)$. By conditional collaboration, $R_{i}(m,\dots,m)=m$, $R_{i}(m+1,\dots,m+1)=m+1$, and $R_{i}(a^{\prime}_{-i})<\max_{j\neq i}a^{\prime}_{j}=m+1$. Hence, $R_{i}(a_{-i})=m$, as desired. ∎
@@
For high-risk public-good games, $R^{*}_{i}$ is welfare-maximizing conditional on safe play.
@@
Let now $a_{i}>R^{*}_{i}(a_{-i})$, that is, $a_{i}$ exceeds the average. Then $\sum_{j\neq i}a_{j}\leq(n-1)a_{i}-1$. For high-risk public-good games, $\lambda<H/(nH-1)$. As the fraction $x/(nx-1)$ is decreasing in $x$, we have $H/(nH-1)\leq a_{i}/(na_{i}-1)$. Furthermore,
@@
    A run is composed of 15 0001500015\ 00015 000 batches of games. Mutation occurs after each batch based on average payoffs within the batch.
@@
    A batch is composed of 10 0001000010\ 00010 000 games. In each game, four reaction functions are drawn at random from the list (with replacement) and payoffs of the highest fixed point are realized. As reactions are monotone, this is easy to compute.
In total, the simulations cover 100×15 000×10 000=15100150001000015100$\times$ 15\ 000$\times$ 10\ $0$ 15 000 × 10 000 = 15 billion games.
@@
0055551010101015151515202020202525252530303030353535354040404045454545505050505555555560606060002222444466668888101010101212121214141414161616161818181820202020Total contribution $\alpha\in\{0,\dots,60\}$ by othersContribution $R_{i}(\alpha)\in\{0,\dots,20\}$
@@
Figure 2 summarizes the 100×500=50 00010050050000100$\times$ 500=50\ $0$ 500 = 50 000 reaction functions that remain at the end of the runs. The solid curve shows the average across these functions. It tracks the average contribution of the others (dashed line) and, importantly, is consistently below it for high $\alpha$. Figure 2 also includes “round down the average” $R^{*}_{i}$ (open circles) and simulation medians (filled dots). These match up remarkably well in the higher end (except at $\alpha=59$).

### v2

@@ 6 Concluding remarks
Blockchain-based smart contracts make contingent commitments both credible and executable. In our framework, players commit to reaction functions and a coordinating contract implements a fixed point of the resulting profile. This is particularly well suited for investment environments, where players can escrow deposits and condition their investment on others’ investments without risking over-exposure. In such settings, reaction functions can help players effectively implement coordination that resolves issues of trust (as in weakest-link games) and disciplines free-riding (as in public-good games). We posit some additional requirements (safe play, norm-proofness, and payoff consistency), which jointly yield simple and practically implementable recommendations.
Compared to the approaches in Tennenholtz (2004) and Kalai et al. (2010), we submit that the reaction functions bring automated play a step closer to practical application. For example, to obtain collaboration in program equilibrium of the Prisoner’s dilemma, Tennenholtz (2004) suggests a program of the form “if my program is identical to my opponent’s program, then the program collaborates; otherwise, it defects”. Conditioning your action on the program of the opponent is fragile, as the smallest difference in programs yields defection. Even though players may be willing to collaborate, programs can be formulated in countless ways and there is a risk that a collaborative outcome would never prevail in practice. For instance, imagine that the other player chooses a different but cooperative-looking program (e.g., a program that unconditionally always collaborates). Even though both are collaborative, the program mismatch still leads to defection. In contrast, our reaction functions specify an action for each action profile of the others. This may be more robust, for instance as illustrated in the cases of weakest-link and public-good games. For example, if players exhibit conditional collaboration in a weakest-link game, we will not automatically resort to the worst-case outcome if some players deviate from the best-reply reaction function.
We end on a few directions for future research. One is to explore reaction functions in games with incomplete information. For instance, the idea of “playing safe” seems appealing even if, say, you do not know the preferences of the other players. A first step in that direction is Theorem 5.2, which shows that the best-reply reaction is weakly dominant against monotone play in high-risk weakest-link games. As weak dominance is a statement only on the player’s own payoffs, knowledge of the others’ payoffs is not important. One can ask more generally, in what games are best-reply reactions $\textit{BR}_{i}$ weakly dominant against monotone play? Theorem 5.2 then points to some conditions: (a) payoff $u_{i}(a)$ is non-decreasing in $a_{-i}\in A_{-i}$, (b) $\textit{BR}_{i}$ is monotone (to ensure fixed points against monotone play), and (c), for each outcome $a\in A$,
@@
There are many games, different from high-risk weakest-link games, for which this applies. The common theme is that (a) $i$ is not harmed by $j$’s actions, that (b) there are some complementarities between player contributions, and (c) that there is a “target” that the player wants to match but not exceed. This target could for instance be derived from Kantian fairness reasoning (e.g. Roemer, 2010; Alger and Weibull, 2013) combined with a strong disutility from being “exploited” (investing more than what is “right” given the investments of the others to the others’ benefit) or some form of inequality aversion (Fehr and Schmidt, 1999). Given the prominent position of the best-reply reaction function, which is simple and intuitive, it would be interesting to map out exactly when it is also optimal.
@@
Consider a potential deviation $R^{\prime}_{i}\in\mathcal{R}_{i}$. If $\mathcal{E}(R^{\prime}_{i},R_{-i})=\emptyset$, then $U_{i}(R^{\prime}_{i},R_{-i})=-\infty<u_{i}(a^{*})$ by convention, so the deviation is not beneficial. For each $a^{\prime}\in\mathcal{E}(R^{\prime}_{i},R_{-i})$ that differs from $a^{*}$, $a^{\prime}_{-i}$ minimizes $u_{i}$ given $a^{\prime}_{i}$. That is,
@@
[…]
@@
That is, $i$ is influenced only by their “neighbors” and takes action $1$ if the player before them chooses 0 whereas the one after them doesn’t. Figure 1 illustrates this case.
ii11i+1i+111i1i-10
@@
As $R_{i}(0,\dots,0)=0$, the outcome $(0,\dots,0)$ is a fixed point. Say player $i$ is unhappy at outcome $a$ if $R_{i}(a_{-i})\neq a_{i}$. We will argue that there are at least two unhappy players at each $a\neq(0,\dots,0)$. For instance, $R_{i}(1,\dots,1)=0$, so all players are unhappy at $(1,\dots,1)$. We restrict to the most challenging cases, namely $a\in\{0,1\}^{n}$. It remains therefore to check outcomes $a$ that include at least one 0 and at least one $1$.
Suppose for contradiction that there is at most one unhappy player at $a$. By construction, $a$ contains at least one sequence of consecutive $1$’s (possibly a single $1$) followed by a 0. The last player $m$ of the sequence is unhappy as $a_{m}=1$ and $a_{m+1}=0$. Therefore, $a$ can contain at most one such sequence (else there would be at least two unhappy players). Moreover, each player $i$ not the first of the sequence is unhappy as $a_{i-1}=a_{i}=1$. Hence, the sequence contains at most two players. There are two options: if $a_{i}=1$ and $a_{j}=0$ otherwise, then i1i-1 and $i$ are unhappy; if $a_{i}=a_{i+1}=1$ and $s_{j}=0$ otherwise, then i1i-1 and $i+1$ are unhappy. (In the second case, that i1i-1 is unhappy hinges on $a_{i-2}=0$. As then $a_{i-2}\neq a_{i+1}$, this uses that $n\geq 4$.) Hence, at all outcomes $a\neq(0,\dots,0)$ there are at least two unhappy players.
@@
If a safe outcome $\underaccent{\bar}{a}\in A$ of the strategic game is a Nash equilibrium, then there is a safe reaction-function equilibrium. In general, if a safe reaction-function equilibrium exists, then there exists a Pareto-efficient safe reaction-function equilibrium.
@@
Part I: Let $R^{*}_{i}\in\mathcal{M}_{i}$ be norm-proof. Fix an amount $\alpha\in\{0,\dots,H\}$ and let $R\in\mathcal{N}$ be such that, for a generic player $i$, $R_{i}(a_{-i})=\alpha$ if $\min_{j\neq i}a_{j}\geq\alpha$ and otherwise $R_{i}(a_{-i})=0$. Then $(\alpha,\dots,\alpha)\in\mathcal{E}(R)$ and $U_{i}(R^{*}_{i},R_{-i})\geq U_{i}(R)$ requires $R^{*}_{i}(\alpha,\dots,\alpha)=\alpha$. As this applies for each $\alpha\in\{0,\dots,H\}$, we conclude that $R^{*}_{i}$ must coordinate if all others do. Next, let $a_{-i}\in A_{-i}$ be such that $\alpha\equiv\min_{j\neq i}a_{j}<\max_{j\neq i}a_{j}\equiv\beta$. As $R^{*}_{i}$ is monotone, $\alpha=R^{*}_{i}(\alpha,\dots,\alpha)\leq R_{i}^{*}(a_{-i})\leq R^{*}_{i}(\beta,\dots,\beta)=\beta$. Hence, $R^{*}_{i}$ exhibits conditional collaboration.
Part II: Let $R^{*}_{i}$ exhibit conditional collaboration. For contradiction, suppose $R^{*}_{i}$ is not norm-proof, so there is $R\in\mathcal{N}$ for which $U_{i}(R^{*}_{i},R_{-i})<U_{i}(R)$. We claim that each $a\in\mathcal{E}(R)$ is coordinated: Without loss, say $a\in\mathcal{E}(R)$ is ordered such that $a_{1}\leq\dots\leq a_{n}$. For contradiction, suppose $a_{1}<a_{n}$. Then $a_{-1}\geqq a_{-n}$. As $R_{i}$ is monotone and $a\in\mathcal{E}(R)$, $a_{1}=R_{1}(a_{-1})\geq R_{n}(a_{-n})=a_{n}$, which is a contradiction. Hence, $U_{i}(R)=u_{i}(\alpha,\dots,\alpha)$ for some $\alpha\in\{0,\dots,H\}$. But $R^{*}_{i}(\alpha,\dots,\alpha)=\alpha$, so $U_{i}(R^{*}_{i},R_{-i})\geq u_{i}(\alpha,\dots,\alpha)=U_{i}(R)$. This is a contradiction. ∎
@@
Let $R_{i}\in\mathcal{R}_{i}$ be norm proof and payoff consistent. Consider an arbitrary $a\in A$ and let $\alpha=\min_{j\neq i}a_{j}$. By payoff consistency, $R_{i}(a_{-i})=R_{i}(\alpha,\dots,\alpha)$. By Observation 3, as $R_{i}$ is norm-proof, $R_{i}(\alpha,\dots,\alpha)=\alpha$. Hence, $R_{i}(a_{-i})=\min_{j\neq i}a_{j}=\textit{BR}_{i}(a_{-i})$. ∎
@@
For public-good games, reaction function $R_{i}\in\mathcal{M}_{i}$ is norm-proof and payoff consistent if and only if there is a monotone function $f_{i}\colon\mathbb{R}\to\mathbb{Z}$ such that, for each $a_{-i}\in A_{-i}$,
$R_{i}(a_{-i})=f_{i}\left(\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right)\in\left\{{\left\lfloor\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rfloor},{\left\lceil\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rceil}\right\}.$
@@
Let $R_{i}\in\mathcal{M}_{i}$ be norm-proof and payoff consistent. Consider an arbitrary $a\in A$ and let
@@
We can “flatten” $a$ to $a^{\prime}\in A$ with $a^{\prime}_{-i}=(m,\dots,m,m+1,\dots,m+1)$ and $\sum_{j\neq i}a^{\prime}_{j}=\sum_{j\neq i}a_{j}$. By payoff consistency, $R_{i}(a_{-i})=R_{i}(a^{\prime}_{-i})$. If $a^{\prime}_{-i}=(m,\dots,m)$, then, by Observation 3 as $R_{i}$ is norm-proof, $R_{i}(a^{\prime}_{-i})=m$ and thus $R_{i}(a_{-i})=m$, as desired. If not, so $a^{\prime}_{-i}\neq(m,\dots,m)$, then $a^{\prime}_{-i}$ contains at least one $m$ and at least one $m+1$. By monotonicity, $R_{i}(m,\dots,m)\leq R_{i}(a^{\prime}_{-i})\leq R_{i}(m+1,\dots,m+1)$. By Observation 3, $R_{i}(m,\dots,m)=m$, $R_{i}(m+1,\dots,m+1)=m+1$, and thus $R_{i}(a^{\prime}_{-i})\in\{m,m+1\}$. Hence, $R_{i}(a_{-i})\in\{m,m+1\}$, as desired. ∎
@@
For high-risk public-good games, $R_{i}\in{\cal M}_{i}$ is welfare-maximizing conditional on safe play if and only if
$R_{i}^{*}(a_{-i})=\left\lfloor\frac{1}{n-1}\textstyle\sum_{j\neq i}a_{j}\right\rfloor.$
@@
Let now $a_{i}>R^{*}_{i}(a_{-i})$, that is, $a_{i}$ exceeds the average. Then $\sum_{j\neq i}a_{j}\leq(n-1)a_{i}-1$. For high-risk public-good games, $\lambda<H/(nH-1)$. As the fraction x/(nx1)x/(nx-1) is decreasing in $x$, we have $H/(nH-1)\leq a_{i}/(na_{i}-1)$. Furthermore,
@@
    A run is composed of 15 00015\ 000 batches of games. Mutation occurs after each batch based on average payoffs within the batch.
@@
    A batch is composed of 10 00010\ 000 games. In each game, four reaction functions are drawn at random from the list (with replacement) and payoffs of the highest fixed point are realized. As reactions are monotone, this is easy to compute.
In total, the simulations cover 100×15 000×10 $\times 15\ 000\times 10\ 0$00=15 billion games.
@@
05510101515202025253030353540404545505055556060022446688101012121414161618182020Total contribution $\alpha\in\{0,\dots,60\}$ by othersContribution $R_{i}(\alpha)\in\{0,\dots,20\}$
@@
Figure 2 summarizes the $100\times 500=50\ 000$ reaction functions that remain at the end of the runs. The solid curve shows the average across these functions. It tracks the average contribution of the others (dashed line) and, importantly, is consistently below it for high $\alpha$. Figure 2 also includes “round down the average” $R^{*}_{i}$ (open circles) and simulation medians (filled dots). These match up remarkably well in the higher end (except at $\alpha=59$).